# CBGBench: Fill in the Blank of Protein-Molecule Complex Binding Graph

### v2

@@ 4 Evaluation Protocol
In previous works, the evaluation usually focuses on two aspects including interaction and chemical property. We extend them and add another two standards on molecule 3D structure’s authenticity with geometry analysis and the 2D graphs reliability with subsutructure analysis, as the following:
Substructure.  To evaluate whether the model succeeds in learning the 2D graph structures for drug molecules. We expand previous metrics to functional groups, as certain functional groups in the drugs act as pharmacophores, such as phenol and pyrimidine. We identify the 25 most frequently occurring functional groups by using EFG [34] (Appendix. A.1 gives the details), and calculated mean frequency of each function group generated in a molecule and the overall multinomial distribution. In the same way, the two values can also be obtained according to rings and atom types. The generated frequency and distribution against the references lead to 6 metrics on substructure analysis including (i) $\bm{\mathrm{JSD}_{\mathrm{at}}}(\downarrow)$, (ii) $\bm{\mathrm{MAE}_{\mathrm{at}}}(\downarrow)$, (iii) $\bm{\mathrm{JSD}_{\mathrm{rt}}}(\downarrow)$, (iv) $\bm{\mathrm{MAE}_{\mathrm{rt}}}(\downarrow)$, (v) $\bm{\mathrm{JSD}_{\mathrm{fg}}}(\downarrow)$, and (vi) $\bm{\mathrm{MEA}_{\mathrm{fg}}}(\downarrow)$, where ‘at’ means atom type, ‘rt’ means ring type and ‘fg’ means functional group.
Chemical Property.  We continue the evaluation of chemical properties from most previous works, including (i) $\bm{\mathrm{QED}}$()($\uparrow$)( ↑ ) as quantitative estimation of drug-likeness; (ii) $\bm{\mathrm{SA}}$()($\uparrow$)( ↑ ) as synthetic accessibility score; (iii) $\bm{\mathrm{LogP}}$ represents the octanol-water partition coefficient, and in general LogP values should be between -0.4 and 5.6 to be good drug candidates [35]; (iv) $\bm{\mathrm{LPSK}}$()($\uparrow$)( ↑ ) as the ratio of the generated drug molecules satisfying the Lipinski’s rule of five.
Table 2: The instance number of training and test split in the datasets for the four tasks and De novo generation. The decomposition is conducted in the training and test set of CrossDocked2020 to avoid label leakage. Side chain decoration and scaffold hopping are two dual tasks, so the training and test set numbers are the same.
@@
Interaction.  In evaluating the generated molecules’ interaction with proteins, the binding affinity is usually the most important metric. Here we unify the affinity calculation with AutoDock Vina for fair comparison. However, previous studies show that larger molecule sizes will lead to a higher probability that the generated molecules can make interaction with the protein, resulting in higher predicted binding affinity [36, 37]. It is also an observation in this paper as the Pearson correlation of Vina Energy v.s. atom number reaches $-0.67$ (See Appendix. A.2). Besides, in some structures, there are more atoms in the protein participating in the interaction, which also causes the average to be unreasonable. Therefore, besides the commonly-used (i) $\bm{E_{\mathrm{vina}}}(\downarrow)$ as the mean Vina Energy, and (ii) $\bm{\mathrm{IMP}\%}(\uparrow)$ as the improvements indicating the ratio of the generated molecule with lower binding energy than the reference, we use mean percent binding gap as (iii) $\bm{\mathrm{MPBG}}(\uparrow)$ ( Appendix. A.3 gives computation) and ligand binding efficacy as (iv) $\bm{\mathrm{LBE}}(\uparrow)$, written as ${\mathrm{LBE}}_{i}=-\frac{{E_{\mathrm{vina}}}}{N_{\mathrm{lig}}}$, indicating how much does a single atom in the molecule contribute to the binding affinity, to eliminate the effects of molecule sizes. Moreover, it is important for the generated molecules to keep the same or similar interaction patterns with proteins [38]. Therefore, we aim to figure out whether the SBDD models can learn the microscopic interaction patterns hidden in the data of the 3D conformations. We use PLIP [39, 40, 41] to characterize the 7 types of interactions, and calculate per-pocket and overall Jensen-Shannon divergence (JSD) between the categorical distributions of generated types and reference types and mean absolute error (MAE) between the frequency of each generated types v.s. it of reference types, in which the frequency is defined as the average number of occurrences of each type of interaction produced by different single generated molecules binding to the same protein, leading to four metrics including (v) $\bm{\mathrm{JSD}_{\mathrm{PP}}}(\downarrow)$, (vi) $\bm{\mathrm{JSD}_{\mathrm{OA}}}(\downarrow)$, (vii) $\bm{\mathrm{MAE}_{\mathrm{PP}}}(\downarrow)$, (viii) $\bm{\mathrm{MAE}_{\mathrm{OA}}}(\downarrow)$, where ‘PP’ and ‘OA’ means per-pocket and overall respectively.
@@
Geometry.  The internal geometry is an important characteristic for distinguishing between general point clouds and molecular structures. The torsion angles [42, 43] are flexible geometries, while bond lengths and bond angles are static to reveal whether the generated molecules have realistic structures. Hence, we evaluate the overall JSD of bond length and angle distribution between reference and generated ones, written as (i) $\bm{\mathrm{JSD}_{\mathrm{BL}}}(\downarrow)$ and (ii) $\bm{\mathrm{JSD}_{\mathrm{BA}}}(\downarrow)$. The other perspective for validating structural resonability is the clash, the occurrence of which is defined when the van der Waals radii overlap by $\geq 0.4$Å [44]. Hence, the ratio of number of atoms generating clashes with protein atoms to the total atom number is written as (ii) $\bm{\mathrm{Ratio}_{\mathrm{cca}}}(\downarrow)$ (cross clashes at atom level). Besides, the ratio of molecules with clashes as $\bm{\mathrm{Ratio}_{\mathrm{cm}}}(\downarrow)$ (molecule with clashes) are evaluated.
@@
The four designation tasks are shown in Figure. 2. We here evaluate the 10 methods in Table. 1, and use Friedman rank [45, 46, 47] as the mean ranking method, to fairly compare the performance of different models in the four aspects. The ranking score is calculated by $(10-\mathrm{rank})$, and the final rank is according to the weighted mean ranking score. Appendix. A.3 gives computation of the metrics.
@@ 5.1.1 Setup
Training.  We use the default configuration in each model’s released code base as the hyper-parameter, and set the training iteration number as 5,000,000 for fair comparison. It is noted that the loss of autoregressive methods exhibits a faster convergence rate of loss, typically requiring only a few tens of thousands of epochs to reach the final best checkpoint. To eliminate the effect brought about by the architecture of GNNs, in implementation, we use GVP [16] and EGNN [15] with GAT [48], as message-passing modules of auto-regressive and diffusion-based models, respectively. Especially, the GNN for encoding and decoding functional groups in D3FG is a combination of GAT and LoCS [49], and CNN for processing atom density maps is four convolution blocks in LiGAN.
Evaluation.  Following the previous protocol, we generate 100 molecules per pocket in the test set, while not all the generated molecules are chemically valid (See Appendix. 19), which will lead to <10,000 molecules. In affinity evaluation, we employ three modes of AutoDock Vina, including ‘Score’, ‘Minimize’, and ‘Dock’. MPBG and LBE are added to the Vina Dock mode. For Vina energy that is larger than 0, we think it is invalid, so we will not report it, and give the lowest ranking in the column of metrics. Moreover, LogP only provides a reference range for drug molecules, so we assigned a rank of 1 to the generated molecules within the range and 2 to those outside the range. In all the tables, values in bold are the best metric, and values with underline are the second and third.
@@ 5.1.2 Result Analysis
LiGAN 0.1167 0.8680 0.3163 0.2701 0.2468 0.0378 4.67
3DSBDD 0.0860 0.8444 0.3188 0.2457 0.2682 0.0494 4.67
GraphBP 0.1642 1.2266 0.5061 0.4382 0.6259 0.0705 9.33
Pocket2Mol 0.0916 1.0497 0.3550 0.3545 0.2961 0.0622 6.67
TargetDiff 0.0533 0.2399 0.2345 0.1559 0.2876 0.0441 2.50
DiffSBDD 0.0529 0.6316 0.3853 0.3437 0.5520 0.0710 6.50
DiffBP 0.2591 1.5491 0.4531 0.4068 0.5346 0.0670 8.83
FLAG 0.1032 1.7665 0.2432 0.3370 0.3634 0.0666 6.83
D3FG 0.0644 0.8154 0.1869 0.2204 0.2511 0.0516 3.33
DecompDiff 0.0431 0.3197 0.2431 0.2006 0.1916 0.0318 1.67
@@
LiGAN 0.46 0.56 0.66 4.39 4.00
3DSBDD 0.48 0.47 0.63 4.72 3.00
GraphBP 0.44 3.29 0.64 4.73 3.25
Pocket2Mol 0.39 2.39 0.65 4.58 4.00
TargetDiff 0.49 1.13 0.60 4.57 3.50
DiffSBDD 0.49 -0.15 0.34 4.89 3.00
DiffBP 0.47 5.27 0.59 4.47 4.50
FLAG 0.41 0.29 0.58 4.93 3.75
D3FG 0.49 1.56 0.66 4.84 1.50
DecompDiff 0.49 1.22 0.66 4.40 3.00
@@
LiGAN -6.47 62.13 -7.14 70.18 -7.70 72.71 4.22 0.3897 0.0346 0.0905 0.1451 0.3416 2.33
3DSBDD  - 3.99 -3.75 17.98 -6.45 31.46 9.18 0.3839 0.0392 0.0934 0.1733 0.4231 5.75
GraphBP  - 0.00  - 1.67 -4.57 10.86 -30.03 0.3200 0.0462 0.1625 0.2101 0.4835 9.33
Pocket2Mol -5.23 31.06 -6.03 38.04 -7.05 48.07 -0.17 0.4115 0.0319 0.2455 0.1535 0.4152 4.16
TargetDiff -5.71 38.21 -6.43 47.09 -7.41 51.99 5.38 0.3537 0.0198 0.0600 0.1757 0.4687 3.08
DiffSBDD  - 12.67 -2.15 22.24 -5.53 29.76 -23.51 0.2920 0.0333 0.1461 0.1777 0.5265 7.42
DiffBP  - 8.60  - 19.68 -7.34 49.24 6.23 0.3481 0.0249 0.1430 0.1256 0.5639 5.67
FLAG  - 0.04  - 3.44 -3.65 11.78 -47.64 0.3319 0.0170 0.0277 0.2762 0.3976 7.50
D3FG  - 3.70 -2.59 11.13 -6.78 28.90 -8.85 0.4009 0.0638 0.0135 0.1850 0.4641 6.67
DecompDiff -5.18 19.66 -6.04 34.84 -7.10 48.31 -1.59 0.3460 0.0215 0.0769 0.1848 0.4369 4.58
@@
Substructure.  Table. 3 gives results on 2D substrucutures (For details, see Appendix. C.1.1). (i) Overall, the auto-regressive models perform worse than the one-shot ones, especially in ring-type evaluation, where the former ones are more likely to generate triangular and tetrahedral rings. (ii) TargetDiff, DiffSBDD and DecompDiff exhibit advantages in atom type. DiffBP has difficulties in generating atoms with low occurrence frequencies. (iii) DecompDiff and D3FG exhibit advantages due to the incorporation of prior knowledge in ring types and functional groups. DiffBP and DiffSBDD perform poorly due to inconsistencies in generating complex fragments.
@@
Chemical Property.  The chemical property is calculated with 2D molecule graphs, so it can be greatly influenced by the molecule substructures. From Table. 4, we can conclude that (i) In terms of the four metrics, the differences among the compared methods are not significant. (ii) D3FG shows the best overall properties, with the highest QED, SA, and LPSK.
@@
Interaction. From Table. 5 shows that (i) LiGAN as a CNN-based method generates molecules that are initialized with high stability according to its superior performance in Vina Score and Vina Min. It also performs well in Vina Dock mode, with a positive MPBG and high LBE, providing good consistency in interaction patterns. (ii) Auto-regressive methods except Pocket2Mol can hardly capture the pocket conditions and generate stably-binding molecules well, with very low IMPin all Vina modes, while Pocket2Mol is the state-of-the-art auto-regressive method in de novo generation. (iii) In diffusion-based methods, TargetDiff outperforms the other competitors in overall interaction comparison, and DecompDiff performs competitively since the difference between these two methods is actually minimal. DiffBP’s performance in docking mode is comparable, but in other modes, it performs less satisfactorily. Performance of DiffSBDD is less than satisfactory. D3FG generates molecules with comparable Vina Energy but small atom numbers, leading to high LBE. For details, see Appendix. C.1.2.

### v3

@@ 4 Evaluation Protocol
In previous works, the evaluation usually focuses on two aspects including interaction and chemical property. While lots of works start to focus another two standards on molecule 3D structure’s authenticity with geometry analysis and the 2D graphs reliability with subsutructure analysis, the evaluation protocol is usually not unified. In this way, we extend them in a unified protocol including the four aspects as the following:
Substructure.  To evaluate whether the model succeeds in learning the 2D graph structures for drug molecules. We expand previous metrics to functional groups, as certain functional groups in the drugs act as pharmacophores, such as phenol and pyrimidine. We identify the 25 most frequently occurring functional groups by using EFG (Salmina et al., 2015) (Appendix. B.1 gives the details), and calculated mean frequency of each function group generated in a molecule and the overall multinomial distribution. In the same way, the two values can also be obtained according to rings and atom types. The generated frequency and distribution against the references lead to 6 metrics on substructure analysis including the Jensen-Shannon divergence as JSD and mean absolute error as MAE, listed as (i) $\bm{\mathrm{JSD}_{\mathrm{at}}}(\downarrow)$, (ii) $\bm{\mathrm{MAE}_{\mathrm{at}}}(\downarrow)$, (iii) $\bm{\mathrm{JSD}_{\mathrm{rt}}}(\downarrow)$, (iv) $\bm{\mathrm{MAE}_{\mathrm{rt}}}(\downarrow)$, (v) $\bm{\mathrm{JSD}_{\mathrm{fg}}}(\downarrow)$, and (vi) $\bm{\mathrm{MAE}_{\mathrm{fg}}}(\downarrow)$, where ‘at’ means atom type, ‘rt’ means ring type and ‘fg’ means functional group.
Chemical Property.  We continue the evaluation of chemical properties from most previous works, including (i) $\bm{\mathrm{QED}}$()($\uparrow$)( ↑ ) as quantitative estimation of drug-likeness; (ii) $\bm{\mathrm{SA}}$()($\uparrow$)( ↑ ) as synthetic accessibility score; (iii) $\bm{\mathrm{LogP}}$ represents the octanol-water partition coefficient, and in general LogP values should be between -0.4 and 5.6 to be good drug candidates (Ghose et al., 1999); (iv) $\bm{\mathrm{LPSK}}$()($\uparrow$)( ↑ ) as the ratio of the generated drug molecules satisfying the Lipinski’s rule of five.
Table 2: The instance number of training and test split in the datasets for the four tasks and De novo generation. The decomposition is conducted separately in the training and test set of CrossDocked2020 to avoid label leakage. Side chain decoration and scaffold hopping are two dual tasks, so the training and test set numbers are the same.
@@
Interaction.  In evaluating the generated molecules’ interaction with proteins, the binding affinity is usually the most important metric. Here we unify the affinity calculation with AutoDock Vina for fair comparison. However, previous studies show that larger molecule sizes will lead to a higher probability that the generated molecules can interact with the protein, resulting in higher predicted binding affinity (Abad-Zapatero & Metz, 2005; Hopkins et al., 2014a). It is also an observation in this paper as the Pearson correlation of Vina Energy v.s. atom number reaches $-0.67$ (See Appendix. B.2). Besides, in some structures, there are more atoms in the protein participating in the interaction, which also causes the average to be unreasonable. Therefore, besides the commonly-used (i) $\bm{E_{\mathrm{vina}}}(\downarrow)$ as the mean Vina Energy, and (ii) $\bm{\mathrm{IMP}\%}(\uparrow)$ as the improvements indicating the ratio of the generated molecule with lower binding energy than the reference, we use mean percent binding gap as (iii) $\bm{\mathrm{MPBG}}(\uparrow)$ (Appendix. B.3 gives computation) and ligand binding efficacy as (iv) $\bm{\mathrm{LBE}}(\uparrow)$, written as ${\mathrm{LBE}}_{i}=-\frac{{E_{\mathrm{vina}}}}{N_{\mathrm{lig}}}$, indicating how much does a single atom in the molecule contribute to the binding affinity, to eliminate the effects of molecule sizes. We give a detailed discussion on the reasonality of LBE in Appendix. B.4. Moreover, it is important for the generated molecules to keep the same or similar interaction patterns with proteins (Zhang et al., 2023a). Therefore, we aim to figure out whether the SBDD models can learn the microscopic interaction patterns hidden in the data of the 3D conformations. We use PLIP (Adasme et al., 2021; li Zuo et al., 2023; Salentin et al., 2015) to characterize the 7 types of interactions and calculate per-pocket and overall JSD between the categorical distributions of generated types and reference types and MAE between the frequency of each generated types v.s. it of reference types, in which the frequency is defined as the average number of occurrences of each type of interaction produced by different single generated molecules binding to the same protein, leading to four metrics including (v) $\bm{\mathrm{JSD}_{\mathrm{PP}}}(\downarrow)$, (vi) $\bm{\mathrm{JSD}_{\mathrm{OA}}}(\downarrow)$, (vii) $\bm{\mathrm{MAE}_{\mathrm{PP}}}(\downarrow)$, (viii) $\bm{\mathrm{MAE}_{\mathrm{OA}}}(\downarrow)$, where ‘PP’ and ‘OA’ means per-pocket and overall respectively.
@@
Geometry.  The internal geometry is an important characteristic for distinguishing between general point clouds and molecular structures. The torsion angles (Jing et al., 2022; Swanson et al., 2023) are flexible geometries, while bond lengths and bond angles are static to reveal whether the generated molecules have realistic structures. Hence, we evaluate the overall JSD of bond length and angle distribution between reference and generated ones, written as (i) $\bm{\mathrm{JSD}_{\mathrm{BL}}}(\downarrow)$ and (ii) $\bm{\mathrm{JSD}_{\mathrm{BA}}}(\downarrow)$. The other perspective for validating structural resonability is the clash, the occurrence of which is defined when the van der Waals radii overlap by $\geq 0.4$Å (Ramachandran et al., 2011). Hence, the ratio of number of atoms generating clashes with protein atoms to the total atom number is written as (iii) $\bm{\mathrm{Ratio}_{\mathrm{cca}}}(\downarrow)$ (cross clashes at atom level). Besides, the ratio of molecules with clashes as (iv) $\bm{\mathrm{Ratio}_{\mathrm{cm}}}(\downarrow)$ (molecule with clashes) are evaluated.
@@
We here evaluate the 12 methods discussed in Table. 1, and use Friedman rank (Friedman, 1940; 1937; Wang et al., 2022) as the mean ranking method, to fairly compare the performance of different models in the four aspects. The ranking score is calculated by $(12-\mathrm{rank})$, and the final rank is according to the weighted mean ranking score. Appendix. B.3 gives detailed computation of the metrics, and their explanations as preliminary. The review of each method’s evaluation aspect in the published papers is given in Apppendix. B.5.
@@ 5.1.1 Setup
Training.  We use the default configuration in each model’s released codebase as the hyper-parameter, and set the training iteration number as 5,000,000 for fair comparison. It is noted that the loss of autoregressive methods exhibits a faster convergence rate of loss, typically requiring only a few tens of thousands of epochs to reach the final best checkpoint. To eliminate the effect brought about by the architecture of GNNs, in implementation, we use GVP (Jing et al., 2020) and EGNN (Satorras et al., 2022) with GAT (Veličković et al., 2018), as message-passing modules of auto-regressive and diffusion-based models, respectively. Especially, the GNN for encoding and decoding functional groups in D3FG is a combination of GAT and LoCS (Kofinas et al., 2022), CNN for processing atom density maps is four convolution blocks in LiGAN, and larger-scaled UNet for VoxBind. The details in hyperparameter tuning and training strategy are given in Appendix. D.
Evaluation.  Following the previous protocol, we generate 100 molecules per pocket in the test set, while not all the generated molecules are chemically valid (See Appendix. 23), which will lead to <10,000 molecules. In affinity evaluation, we employ three modes of AutoDock Vina, including ‘Score’, ‘Minimize’, and ‘Dock’. MPBG and LBE are added to the Vina Dock mode. For Vina energy that is larger than 0, we think it is invalid, so we will not report it, and give the lowest ranking in the column of metrics. Moreover, LogP only provides a reference range for drug molecules, so we assigned a rank of 1 to the generated molecules within the rank of 2 to those outside the range. In all the tables, values in bold are the best metric, and values with underline are the second and third.
@@ 5.1.2 Result Analysis
LiGAN 0.1167 0.8680 0.3163 0.2701 0.2468 0.0378 6.33
3DSBDD 0.0860 0.8444 0.3188 0.2457 0.2682 0.0494 6.33
GraphBP 0.1642 1.2266 0.5061 0.4382 0.6259 0.0705 11.17
Pocket2Mol 0.0916 1.0497 0.3550 0.3545 0.2961 0.0622 8.33
TargetDiff 0.0533 0.2399 0.2345 0.1559 0.2876 0.0441 3.33
DiffSBDD 0.0529 0.6316 0.3853 0.3437 0.5520 0.0710 8.16
DiffBP 0.2591 1.5491 0.4531 0.4068 0.5346 0.0670 10.67
FLAG 0.1032 1.7665 0.2432 0.3370 0.3634 0.0666 8.50
D3FG 0.0644 0.8154 0.1869 0.2204 0.2511 0.0516 4.67
DecompDiff 0.0431 0.3197 0.2431 0.2006 0.1916 0.0318 2.50
MolCraft 0.0490 0.3208 0.2469 0.0264 0.1196 0.0477 3.00
VoxBind 0.0942 0.3564 0.2401 0.0301 0.1053 0.0761 5.00
@@
LiGAN 0.46 0.56 0.66 4.39 4.75
3DSBDD 0.48 0.47 0.63 4.72 3.50
GraphBP 0.44 3.29 0.64 4.73 3.75
Pocket2Mol 0.39 2.39 0.65 4.58 4.75
TargetDiff 0.49 1.13 0.60 4.57 4.25
DiffSBDD 0.49 -0.15 0.34 4.89 3.50
DiffBP 0.47 5.27 0.59 4.47 5.25
FLAG 0.41 0.29 0.58 4.93 4.25
D3FG 0.49 1.56 0.66 4.84 2.00
DecompDiff 0.49 1.22 0.66 4.40 3.75
MolCraft 0.48 0.87 0.66 4.39 4.00
VoxBind 0.54 2.22 0.65 4.70 2.75
@@
LiGAN -6.47 62.13 -7.14 70.18 -7.70 72.71 4.22 0.3897 0.0346 0.0905 0.1451 0.3416 2.91
3DSBDD  - 3.99 -3.75 17.98 -6.45 31.46 9.18 0.3839 0.0392 0.0934 0.1733 0.4231 7.17
GraphBP  - 0.00  - 1.67 -4.57 10.86 -30.03 0.3200 0.0462 0.1625 0.2101 0.4835 11.33
Pocket2Mol -5.23 31.06 -6.03 38.04 -7.05 48.07 -0.17 0.4115 0.0319 0.2455 0.1535 0.4152 5.67
TargetDiff -5.71 38.21 -6.43 47.09 -7.41 51.99 5.38 0.3537 0.0198 0.0600 0.1757 0.4687 4.67
DiffSBDD  - 12.67 -2.15 22.24 -5.53 29.76 -23.51 0.2920 0.0333 0.1461 0.1777 0.5265 9.25
DiffBP  - 8.60  - 19.68 -7.34 49.24 6.23 0.3481 0.0249 0.1430 0.1256 0.5639 7.41
FLAG  - 0.04  - 3.44 -3.65 11.78 -47.64 0.3319 0.0170 0.0277 0.2762 0.3976 9.00
D3FG  - 3.70 -2.59 11.13 -6.78 28.90 -8.85 0.4009 0.0638 0.0135 0.1850 0.4641 8.17
DecompDiff -5.18 19.66 -6.04 34.84 -7.10 48.31 -1.59 0.3460 0.0215 0.0769 0.1848 0.4369 6.08
MolCraft -6.15 54.25 -6.99 56.43 -7.79 56.22 8.38 0.3638 0.0214 0.0780 0.1868 0.4574 3.75
VoxBind -6.16 41.80 -6.82 50.02 -7.68 52.91 9.89 0.3588 0.0257 0.0533 0.1850 0.4606 4.00
@@
Substructure.  Table. 3 gives results on 2D substructures, showing that (i) Overall, the auto-regressive models perform worse than the one-shot ones, especially in ring-type evaluation, where the former ones are more likely to generate triangular and tetrahedral rings. (ii) Diffusion-based models also exhibit advantages in atom type, according to the results of TargetDiff and MolCraft, except for DiffBP has difficulties in generating atoms with low occurrence frequencies. (iii) Besides VoxBind, MolCraft shows the overall superiority in complex functional group generation, as it states in addressing the mode collapse issues, while DecompDiff and D3FG also exhibit advantages due to the incorporation of prior knowledge in ring types and functional groups. DiffBP and DiffSBDD perform poorly due to inconsistencies in generating complex fragments. For a detailed comparison, see Appendix. E.1.1
@@
Chemical Property.  The chemical property is calculated with 2D molecule graphs, so it can be greatly influenced by the molecule substructures. From Table. 4, we can conclude that (i) In terms of the four metrics, the differences among the compared methods are not significant. (ii) D3FG shows the best overall properties, with the competitive QED, SA, and LPSK.
@@
Interaction. From Table. 5 shows that (i) LiGAN and VoxBind as CNN-based methods generate molecules that are initialized with high stability according to its competitive performance in Vina Score and Vina Min. It also performs well in Vina Dock mode, with a positive MPBG and high LBE, providing good consistency in interaction patterns. (ii) Auto-regressive methods except Pocket2Mol can hardly capture the pocket conditions and generate stably-binding molecules well, with very low IMPin all Vina modes, while Pocket2Mol is the state-of-the-art auto-regressive method in de novo generation. (iii) In diffusion-based methods, MolCraft outperform the other competitors in overall interaction comparison, and TargetDiff and DecompDiff perform competitively since the difference between these two methods is minimal. DiffBP’s performance in docking mode is comparable, but in other modes, it performs less satisfactorily. Performance of DiffSBDD is less than satisfactory. D3FG generates molecules with comparable Vina Energy but small atom numbers, leading to high LBE. For details, see Appendix. E.1.2.
Geometry.  From Table. 6, conclusions can be drawn that (i) MolCraft generates overall the most realistic structures, according to the internal geometries, and DecompDiff is comparable. (ii) The molecule-protein clashes can usually be avoided by diffusion-based models. However, in the auto-regressive models, only Pocket2Mol avoids clashes. The high frequency of clashes in auto-regressive methods can be explained as follows: These models first identify frontier amino acids or atoms within the molecule. During the placing of new atoms, incorrect binding site localization leads to an erroneous direction for the molecule’s auto-regressively growing path, causing it to extend inward towards the protein. For detailed results, see Appendix. E.1.3.
Conclusion and Discussion.
From Table. 7 and the previous discussion, we conclude that
  • (i)
    The CNN-based methods like LiGAN and VoxBind are highly competitive, especially in aspects of Interaction, which explains why such methods in the field of drug design and molecule generation still prevail in recent years. This is partly attributed to the fact that CNNs have an advantage over GNNs in perceiving many-body patterns within a single filter (Townshend et al., 2021). As a result, it encourages further research into GNNs for 3D point clouds to develop architectures that can match the expressivity of CNNs.
  • (ii)
    MolCraft achieves the best overall performance, with TargetDiff closely following. In contrast, DecompDiff and D3FG as the variants of TargetDiff that incorporate domain knowledge show some degeneration in performance. This observation reveals that the current incorporation of physicochemical priors can hardly improve the quality of generated molecules. Effectively integrating domain knowledge to guide the model to generate structurally sound molecules remains a challenge. For example, atom clashes are very common in generated molecules. While DecompDiff employs the prior guidance for this, the problem has still not been fully solved. Integrating domain knowledge into the training process may mitigate this issue (Huang et al., 2024a; Adams & Coley, 2023).
  • (iii)
    Only Pocket2Mol as an auto-regressive method achieves competitive results, which we attribute to the following reasons: First, it utilizes chemical bonds to constrain atoms to grow orderly along chemical bonds rather than to grow based on distance to the pocket as in DiffBP. Second, it simultaneously predicts the bond types and employs contrastive learning by sampling positive and negative instances of atom positions as real and fake bonds, which is not fully considered by FLAG, enhancing the model’s ability to perceive chemical bond patterns. Therefore, we believe that enabling the autoregressive methods model to successfully capture the patterns of chemical bonds is very essential.
5.2 Extension on Subtasks
Setup.  Lead optimization is to strengthen the function or property of the binding molecules by remodeling the existing partial 3D graph of molecules. We show the interaction analysis and chemical property in the main context and omit the interaction pattern analysis since maintaining the patterns is not necessary for lead optimization. The domain-knowledge-based methods can hardly be transferred for these tasks since different tasks require different priors, so we have not compared them here. Besides, the voxelize-based methods are not easily extended to these tasks, and we regard the transferring of methods like LiGAN and 3DSBDD as future work. Hence, we here compare 6 methods that model atoms’ continuous positions with GNNs and have not employed domain knowledge. When training, since the number of training instances is smaller, we use the pretrained models on de novo generation and finetune the auto-regressive models with 1,000,000 iterations. For diffusion-based models with one-shot generation, we train them from scratch because the zero-center-of-mass (Satorras et al., 2021) technique is shifted from employing protein geometric centers to using molecule context’s ones. For detailed results, please refer to Appendix. E.2.